# FoilDiff: A Hybrid Transformer Backbone for Diffusion-based Modelling of 2D Airfoil Flow Fields

### v1

@@ Abstract
The accurate prediction of flow fields around airfoils is crucial for aerodynamic design and optimisation. Computational Fluid Dynamics (CFD) models are effective but computationally expensive, thus inspiring the development of surrogate models to enable quicker predictions. These surrogate models can be based on deep learning architectures, such as Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), and Diffusion Models (DMs). Diffusion models have shown significant promise in predicting complex flow fields. In this work, we propose FoilDiff, a diffusion-based surrogate model with a hybrid-backbone denoising network. This hybrid design combines the power of convolutional feature extraction and transformer-based global attention to generate more adaptable and accurate representations of flow structures. FoilDiff takes advantage of Denoising Diffusion Implicit Model (DDIM) sampling to optimise the efficiency of the sampling process at no additional cost to model generalisation. We used encoded representations of Reynolds number, angle of attack, and airfoil geometry to define the input space for generalisation across a wide range of aerodynamic conditions. When evaluated against state-of-the-art models, FoilDiff shows significant performance improvements, with mean prediction errors reducing by up to 85% on the same datasets. The results have demonstrated that FoilDiff can provide both more accurate predictions and better-calibrated predictive uncertainty than existing diffusion-based models.
@@ 1 Introduction
The accurate and efficient prediction of flow fields across airfoils is crucial for advancing aerodynamic design methodologies of many engineering systems. Computational Fluid Dynamics (CFD) techniques, particularly those based on Reynolds-averaged Navier-Stokes (RANS) equations, are currently used to model flow behaviour [1]. However, the numerical complexity of these equations necessitate computationally expensive simulations, making them a bottleneck in iterative design and optimization processes. As a result, surrogate modelling approaches have gained traction as an effective means to balance computational efficiency and predictive accuracy. Deep learning approaches have demonstrated significant promise in approximating RANS solutions across a wide range of flow conditions [2, 3, 4]. Diffusion models are a class of generative models that have been used more recently in flow field predictions, as an alternative to more established deep learning surrogate modelling methodologies, such as Convolutional Neural Networks (CNNs) and standalone U-Net models [5, 6, 7].
@@
CNNs have played significant roles in flow field prediction due to their ability to efficiently extract spatial features [8, 9, 10]. CNNs have been used for real-time modelling of non-uniform steady laminar flow, achieving velocity predictions four orders of magnitude faster than CFD solvers with a minor trade off in accuracy [4]. Similarly, they have also been utilized to extract geometric parameters from airfoil shapes, which along with the Reynolds number and angle of attack, were fed into a Multilayer Perceptron (MLP) to approximate flow fields [10]. Advancements with encoder-decoder architectures, such as the CNN-FOIL [11], have enabled efficient transonic flow field predictions by enhancing shared feature extraction across multiple aerodynamic parameters. Furthermore, the integration of attention mechanisms and U-Net architectures has significantly improved information capture and predictive accuracy in airfoil flow field models [12, 13, 14].
The growing success of deep learning in fluid dynamics has been further propelled by the rise of generative models, which have achieved remarkable results in various applications, including image synthesis [15], material design [16], and turbulence generation [17]. In the context of CFD, Generative Adversarial Networks (GANs) models have been applied to turbulence field resolution [18] and flow field prediction using Variational Autoencoders (VAEs) [19]. Diffusion models, in particular, have emerged as a powerful alternative offering stable training and enhanced mode coverage compared to GANs [20, 21, 22].
@@
Recent studies have demonstrated the effectiveness of denoising diffusion models over standalone CNNs and U-Nets in capturing complex fluid dynamics [5, 23, 6, 7]. Currently, diffusion models applied to the task of airfoil flow field prediction rely on attention augmented U-Nets or standard diffusion transformers as the model backbones [7, 24]. Each approach comes with certain strengths and limitations. U-Nets are effective at capturing local dependencies and features but struggle with long-range dependencies [25]. Attention-augmented U-Nets omit the feed-forward sublayers that are fundamental to standard Transformer blocks, rather using the attention gates to compute attention coefficients by comparing encoder feature maps with gating signals from the decoder [26]. As a result, while they are capable of identifying where long-range interactions should occur, they often lack depth necessary to encode these dependencies into globally coherent representations of fluid dynamics. In contrast, transformer-based diffusion models, such as Diffusion Transformers (DiT) [27], have demonstrated improved generalization capabilities, relying on the global modelling ability of Vision Transformers (ViTs) [28], to yield high fidelity predictions. DiTs introduce separated encoders and decoders which are often separately trained to encode inputs to latent space, which stand to benefit from U-Net-style connections at each resolution layer [29].
To address these limitations, we introduce FoilDiff. The FoilDiff backbone combines the benefits of a connected encoder-decoder with a latent transformer. Additionally, we provide deep conditioning to the latent bottleneck representation to enable more accurate modelling. Finally, FoilDiff employs the Denoising Diffusion Implicit Model (DDIM) accelerated sampling strategy to enhance computational efficiency while maintaining predictive accuracy [30].
@@
To evaluate FoilDiff, we assess its ability to predict steady-state flow fields around the RAF30 airfoil across a range of Reynolds numbers and angles of attack. We compare performance against state-of-the-art diffusion models with U-Net and attention-augmented U-Net backbones. FoilDiff outperforms these baselines in both predictive accuracy and uncertainty modelling. We further conduct ablation studies to isolate the contributions of each architectural enhancement, highlighting their role in improving model efficiency and expressiveness.
@@
The rest of this article is structured as follows: Section 2 provides an overview of diffusion models, detailing the underlying architecture and enhanced sampling strategies. Section 3 outlines the proposed model architecture, training dataset and configurations. Section 4 presents experimental results, comparing the performance of our approach to existing models. Finally, Section 5 concludes the paper with key findings and potential directions for future research.
@@ 2 Preliminary Background on Diffusion Models
Diffusion models are inspired by the principles of non-equilibrium thermodynamics [31]. These models operate in a two-process framework: a forward process where noise is added to the data in a gradual process, and a reverse process, in which the model learns how to generate the original data from the noisy input.
@@ 2.1 Denoising Diffusion Probabilistic Models (DDPMs)
In DDPMs [32], the objective, as depicted in Figure 1, is to learn the conditional distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, an estimate of an ideal $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, which reverses the forward process $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$.
@@
In the forward process, noise is progressively added to the data vector $\mathbf{x}_{0}$ across a number of time steps $T$ , transitioning from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$ to $\dots$ to $\mathbf{x}_{T}$. This progression follows a Gaussian conditional distribution defined as:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}I),$ (2.1)
where $\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}$ is the mean, $\beta_{t}I$ represents the covariance matrix, $I$ is the identity matrix, and $\beta_{t}\in(0,1)$ is the noise level parameter set according to a predefined schedule such that $\beta_{1}<\beta_{2}<\dots<\beta_{T}$.
As the forward process proceeds, the parameter $\beta_{t}$ increases, causing the mean $\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}$ to drift farther from the previous state $\mathbf{x}_{t-1}$ and the variance $\beta_{t}I$ to increase. Thus, as $t$ approaches $T$, the data $\mathbf{x}_{t}$ becomes increasingly noisy. This process is represented as a Markov chain [33], adhering to the property that each state depends only on its immediate predecessor:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{x}_{t-2},\dots,\mathbf{x}_{0})=q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ (2.2)
Therefore, the joint distribution across all states in the forward process can be expressed in the form:
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),$ (2.3)
where $\mathbf{x}_{1:T}$ represents all intermediate states from $\mathbf{x}_{1}$ to $\mathbf{x}_{T}$, and $\mathbf{x}_{0}$ is the original data. The state $\mathbf{x}_{t}$ can then be expressed as:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon}$ (2.4)
a linear combination of $\mathbf{x}_{0}$ and a noise variable $\mathbf{\epsilon}$, where $\alpha_{t}$ denotes the cumulative product.
$\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ (2.5)
Sampling begins with a noisy input and iteratively generates less noisy samples, continuing until $\mathbf{x}_{0}$ is reached from $\mathbf{x}_{T}$.
$p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod_{t=1}^{T}p(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ (2.6)
To achieve this goal, a network with parameters $\theta$ is used to model each step of the reverse process,
@@
where $\mu_{\theta}(\mathbf{x}_{t};t)\in\mathbb{R}^{d}$ and $\Sigma_{\theta}(\mathbf{x}_{t};t)\in\mathbb{R}^{d\times d}$ are the mean and covariance of the distribution at step $t$, respectively [34]. The forward process only requires adding noise to the data; it does not involve any learning and remains fixed. As a result, only the reverse process needs to be trained in diffusion models. This distinguishes diffusion models from VAEs [35, 36].
The objective remains to learn the conditional distribution, $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, an estimate of the true posterior $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, which reverses the forward diffusion process $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. Since the exact posterior is intractable, the model is trained to minimize the Kullback-Leibler (KL) divergence between the learned and true distributions, ensuring that the reverse process closely follows the true generative trajectory. The training objective is given by:
$\mathcal{L}_{\text{KL}}=\mathbb{E}_{q(\mathbf{x}_{0:T})}\sum_{t=1}^{T}D_{\text{KL}}(q(\mathbf{x}_{t-1}|\mathbf{x}_{t})\parallel p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})),$ (2.8)
where $D_{\text{KL}}(q\parallel p)$ denotes the KL divergence between the true and approximated distributions. The training of this generator model focuses on its ability to iteratively produce a denoised output $\mathbf{x}_{t-1}$ from a given noisy input $\mathbf{x}_{t}$. Following formulation in Equation 2.4, the prediction for $\mathbf{x}_{t-1}$ based on $\mathbf{x}_{t}$ can be expressed as a function of $\mathbf{\epsilon}_{\theta}$, the output of the noise predictor neural network parameterised by $\theta$.
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\mathbf{\epsilon}_{\theta})$ (2.9)
This reverse process is integral to reconstructing the original data distribution from noise. In the canonical process for training these models effectively [32], a minibatch is created by randomly sampling a data point $\mathbf{x}_{0}$ from the dataset, selecting a time step $t$, and introducing noise $\mathbf{\epsilon}$ drawn from a standard normal distribution $\mathcal{N}(0,I)$. These steps collectively produce a noisy sample $\mathbf{x}_{t}$ at the chosen time step, as visualized in Figure 3. The training objective is then to minimize the discrepancy between the actual noise $\mathbf{\epsilon}$ and the noise predicted by the model, represented as $\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)$. This objective is formalized using the loss function:
$L_{\theta}=\mathbb{E}_{\mathbf{x},\mathbf{\epsilon}\sim\mathcal{N}(0,I)}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right]$ (2.10)
By minimizing this loss function, the model, parameterised by $\theta$, learns to accurately estimate the noise component within a noisy sample $\mathbf{x}_{t}$. This training process is pivotal as it equips the model with the capability to progressively reduce noise and approach the true data distribution during the reverse diffusion process. The training framework reinforces the model’s denoising capability, which is essential for generating high-quality samples in the reverse process. This step-by-step denoising operation ensures that the model not only learns to replicate the distribution of the original data but also excels at generating new samples that closely align with the target distribution. Consequently, diffusion models trained under this paradigm serve as powerful generative models capable of producing realistic and coherent data samples.

### v2

# FoilDiff: A Hybrid Diffusion Transformer Model for Airfoil Flow Field Prediction

@@ Abstract
The accurate prediction of flow fields around airfoils is crucial for aerodynamic design and optimisation. Computational Fluid Dynamics (CFD) models are effective but computationally expensive. This has inspired the development of surrogate models to enable quicker predictions. These surrogate models can be based on deep learning architectures, such as Convolutional Neural Networks (CNNs), Graph Neural Networks (GNNs), and Diffusion Models (DMs). Diffusion models have shown significant promise in predicting complex flow fields. In this work, we propose FoilDiff, a diffusion-based surrogate model with a hybrid-backbone denoising network. This hybrid design combines the power of convolutional feature extraction and transformer-based global attention to generate more adaptable and accurate representations of flow structures. FoilDiff takes advantage of Denoising Diffusion Implicit Model (DDIM) sampling to optimise the efficiency of the sampling process at no additional cost to model generalisation. We use encoded representations of Reynolds number, angle of attack, and airfoil geometry to define the input space for generalisation across a wide range of aerodynamic conditions. When evaluated against state-of-the-art models, FoilDiff shows significant performance improvements, with mean prediction errors reduced by up to 85% on the same datasets. These results have demonstrated that FoilDiff can provide both more accurate predictions and better-calibrated predictive uncertainty than existing diffusion-based models.
@@ 1 Introduction
The accurate and efficient prediction of flow fields across airfoils is crucial for advancing aerodynamic design methodologies of many engineering systems. Computational Fluid Dynamics (CFD) techniques, particularly those based on Reynolds-averaged Navier-Stokes (RANS) equations, are currently used to model flow behaviour [5MATYUSHENKO201715]. However, the numerical complexity of these equations necessitates computationally expensive simulations, making them a bottleneck in iterative design and optimisation processes. As a result, surrogate modelling approaches have gained traction as an effective means to balance computational efficiency and predictive accuracy. Deep learning approaches have demonstrated significant promise in approximating RANS solutions across a wide range of flow conditions [Bruni2023, Thuerey2018, GuoCNN]. Diffusion models are a class of generative models that have been used more recently in flow field predictions, as an alternative to more established deep learning surrogate modelling methodologies, such as Convolutional Neural Networks (CNNs) and standalone U-Net models [gefan2023denoisingdiffusionmodelfluid, hu2024generativepredictionflowfield, Liu2023]. Diffusion models for airfoil flow field prediction, while demonstrating strong generative capability, often require exceptionally lengthy training runs in the tens of millions of iterations or the training of additional models [Liu2023, xiang2024aeroditdiffusiontransformersreynoldsaveraged]. This substantial computational demand presents a major bottleneck, underscoring the need for more efficient training strategies or model architectures to make diffusion-based surrogates more practically viable for aerodynamic design workflows.
@@
CNNs have played significant roles in flow field prediction due to their ability to efficiently extract spatial features [wang2024recentadvancesmachinelearning, ronneberger2015unetconvolutionalnetworksbiomedical, SekarCNN]. CNNs have been used for real-time modelling of non-uniform steady laminar flow, achieving velocity predictions four orders of magnitude faster than CFD solvers with a minor trade-off in accuracy [GuoCNN]. Similarly, they have also been utilized to extract geometric parameters from airfoil shapes, which along with the Reynolds number and angle of attack, were fed into a Multilayer Perceptron (MLP) to approximate flow fields [SekarCNN]. Advancements with encoder-decoder architectures, such as the CNN-FOIL [Duru2022], have enabled efficient transonic flow field predictions by enhancing shared feature extraction across multiple aerodynamic parameters. Furthermore, the integration of attention mechanisms and U-Net architectures has significantly improved information capture and predictive accuracy in airfoil flow field models [Wang2024, Zuo2024, Bruni2024].
The growing success of deep learning in fluid dynamics has been further propelled by the rise of generative models, which have achieved remarkable results in various applications, including image synthesis [dhariwal2021diffusionmodelsbeatgans], material design [fu2022materialstransformerslanguagemodels], and turbulence generation [gao2023bayesianconditionaldiffusionmodels]. In the context of CFD, Generative Adversarial Networks (GAN) models have been applied to turbulence field resolution [kimpredictioncontroltwodimensionaldecaying] and flow field prediction using Variational Autoencoders (VAEs) [ranade2021composableautoencoderbasediterativealgorithm]. An early VAE model for airfoil flow field prediction [Wang2021], learned a compact mapping from airfoil geometry and flow conditions to full-field pressure and velocity distributions. The model reproduced high quality results at a fraction of the computational cost but tended to smooth sharp gradients near shocks and trailing edges, a common limitation of VAEs with Gaussian latent spaces. To improve generalisation under off-design conditions, a prior Variational Autoencoder (pVAE) was proposed [Yang2022]. This incorporated a cruise-condition flow field as a latent prior and introduced physics-based loss terms to enforce aerodynamic force balance and mass conservation. To extend this, GAN models adopted an adversarial approach through a data-augmented conditional GAN (daGAN) combining a U-Net-based generator and a discriminator to sharpen flow reconstructions [WU2022]. Data augmentation improved robustness, and the adversarial objective mitigated over-smoothing, yielding clearer shocks and wake features. Despite its success, daGAN still faced typical GAN challenges such as training instability. Diffusion models have emerged as a powerful alternative offering stable training and enhanced mode coverage compared to GANs [liu2021fasterstabilizedgantraining, abaidi2024exploring, dhariwal2021diffusionmodelsbeatgans, liu2023combatingmodecollapsegans].
@@
Recent studies have demonstrated the effectiveness of denoising diffusion models over standalone CNNs and U-Nets in capturing complex fluid dynamics [gefan2023denoisingdiffusionmodelfluid, SHU2023111972, hu2024generativepredictionflowfield, Liu2023]. Currently, diffusion models applied to the task of airfoil flow field prediction rely on attention augmented U-Nets or standard diffusion transformers as the model backbones [Liu2023, xiang2024aeroditdiffusiontransformersreynoldsaveraged]. Each approach comes with certain strengths and limitations. U-Nets are effective at capturing local dependencies and features but struggle with long-range dependencies [rombach2022highresolutionimagesynthesislatent]. Attention-augmented U-Nets omit the feed-forward sublayers that are fundamental to standard Transformer blocks, rather using the attention gates to compute attention coefficients by comparing encoder feature maps with gating signals from the decoder [Oktay2018AttentionUNet]. As a result, while they are capable of identifying where long-range interactions should occur, they often lack the depth necessary to encode these dependencies into globally coherent representations of fluid dynamics. This is the architecture applied by the state-of-the-art model Aifnet [Liu2023]. This model demonstrated superior accuracy over standalone UNets in modelling the epistemic uncertainty present in CFD simulations of highly turbulent flows. Despite its strong predictive accuracy, the model’s reliance on millions of training iterations highlights inefficiencies in its learning process, and its generalisation capability remains suboptimal. In contrast, transformer-based diffusion models, such as Diffusion Transformers (DiT) [peebles2023scalablediffusionmodelstransformers], have demonstrated improved generalisation capabilities over UNet-based diffusion models, relying on the global modelling ability of Vision Transformers (ViTs) [dosovitskiy2021image], to yield high-fidelity predictions. DiTs for airfoil flow field prediction such as AeroDiT [xiang2024aeroditdiffusiontransformersreynoldsaveraged] introduce separated encoders and decoders to encode inputs into latent space, this approach involves the training of three models, two VAEs and one transformer all trained separately with the trained VAEs used in the transformer training. This multi-stage training process introduces additional computational overhead and tuning complexity. Additionally, the separated encoders and decoders stand to benefit from U-Net-style connections at each resolution layer, to provide additional control and fine tuning over the encoding and decoding to and from the latent space [VAEskips].
@@
To address these limitations, we introduce FoilDiff. FoilDiff combines the benefits of a connected encoder-decoder with a latent transformer within one training process. Additionally, we provide deep conditioning to the latent bottleneck representation to enable more accurate modelling. Finally, FoilDiff employs the Denoising Diffusion Implicit Model (DDIM) accelerated sampling strategy to enhance computational efficiency while maintaining predictive accuracy [song2022denoisingdiffusionimplicitmodels]. To evaluate FoilDiff, we assess its ability to predict steady-state flow fields around across a range of airfoils, Reynolds numbers and angles of attack. We compare performance against state-of-the-art diffusion models with U-Net and transformer backbones. FoilDiff outperforms these baselines in both predictive accuracy and uncertainty modelling. As some state-of-the-art models differ substantially in scale, rendering direct comparison inappropriate, we further conduct ablation studies to isolate the contributions of each architectural enhancement, highlighting their role in improving model efficiency and expressiveness.
@@
The rest of this article is structured as follows: Section 2 provides an overview of diffusion models, detailing the underlying architecture and enhanced sampling strategies. Section 3 outlines the proposed model architecture, training and inference configurations. Section 4 presents experimental analysis, comparing the performance of our approach to existing models. Finally, Section 5 concludes the paper with key findings and potential directions for future research.
@@ 2 Preliminary Background on Diffusion Models
Diffusion models are inspired by the principles of non-equilibrium thermodynamics [pmlr-v37-sohl-dickstein15]. These models operate in a two-process framework: a forward process where noise is added to the data in a gradual process, and a reverse process, in which the model learns how to generate the original data from the noisy input.
@@ 2.1 Denoising Diffusion Probabilistic Models (DDPMs)
DDPMs learn to reverse the forward process $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ with the conditional distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, an estimate of $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ [Ho2020].
@@
In the forward process, noise is progressively added to the data vector $\mathbf{x}_{0}$ across a number of time steps $T$ , transitioning from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$ to $\dots$ to $\mathbf{x}_{T}$. The intermediate states $\mathbf{x}_{t}$ can then be expressed as:
$\mathbf{x}_{t}=\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon}$ (2.1)
a linear combination of $\mathbf{x}_{0}$ and a noise variable $\boldsymbol{\epsilon}$, where $\overline{\alpha}_{t}$ denotes the cumulative product of the variance schedule [Ho2020]
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\overline{\alpha}_{t}}{x_{0}},(1-\overline{\alpha}_{t})I),$ (2.2)
The reverse process then begins with a noisy input and iteratively generates less noisy samples, continuing until $\mathbf{x}_{0}$ is reached from $\mathbf{x}_{T}$. To achieve this goal, a network with parameters $\theta$ is used to model each step of the reverse process,
@@
where $\mu_{\theta}(\mathbf{x}_{t};t)\in\mathbb{R}^{d}$ and $\Sigma_{\theta}(\mathbf{x}_{t};t)\in\mathbb{R}^{d\times d}$ are the mean and covariance of the distribution at step $t$, respectively [ghojogh_ghodsi_2024]. The forward process only requires adding noise to the data; it does not involve any learning and remains fixed. As a result, only the reverse process needs to be trained in diffusion models. This distinguishes diffusion models from VAEs [kingma2022autoencodingvariationalbayes, inbookGhoj]. Following formulation in Equation 2.1, the prediction for $\mathbf{x}_{t-1}$ based on $\mathbf{x}_{t}$ can be expressed as a function of $\boldsymbol{\epsilon}_{\theta}$, the output of the noise predictor neural network parameterised by $\theta$.
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\mathbf{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon}_{\theta})$ (2.4)
The training objective is then to minimise the discrepancy between the actual noise $\boldsymbol{\epsilon}$ and the noise predicted by the model, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$. This objective is formalized using the loss function:
$L_{\theta}=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)}\left[\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right]$ (2.5)
By minimizing this loss function, the model, parameterised by $\theta$, learns to accurately estimate the noise component within a noisy sample $\mathbf{x}_{t}$. Consequently, diffusion models trained under this paradigm serve as powerful generative models capable of producing realistic and coherent data samples.